# Stochastic and incremental subgradient methods for convex optimization on Hadamard spaces

### v3

@@ 6 Subgradient-based minimization of sums
We denote the projection onto a nonempty closed geodesically convex set $C$ by $P_{C}$ (well-defined by [13, Proposition II.2.4]). Define an oracle Busemann that accepts a Busemann subdifferentiable function $g\colon C\to\mathbb{R}$ and a point $x\in C$, and returns a Busemann subgradient for $g$ at $x$: $\xi,s]\leftarrow\texttt{Busemann}(g,x)$. If $\xi,s]$ is a Busemann subgradient for $g$ at $x$, we will use the notation $r_{x,\xi}$ to denote the ray issuing from $x$ with direction $\xi$.
@@
    The objective $f\colon C\to\mathbb{R}$ decomposes as $f=\sum_{i=1}^{m}f_{i}$, where each $f_{i}\colon C\to\mathbb{R},i=1,\dots,m$ is Busemann subdifferentiable.
@@
$d(x^{+},y)^{2}$ $=d\left(P_{C}(r\left(st\right)),P_{C}(y)\right)^{2}$
$\leq d\left(r\left(st\right),y\right)^{2}$
@@
The second line uses nonexpansivity of the projection $P_{C}$. Letting $\delta\to\infty$ and using once more (5) we deduce
@@
The analysis of the stochastic subgradient method borrows from [5, Lemma 3.6] and [7, Theorem 8.35]. Define the notation $f^{k}_{\text{best}}:=\min_{i=1,\dots,k}f(x^{i})-f_{\text{opt}}$. For stochastic terminology, see [34].
@@
Finally, since the diameter of $C$ is at most $D$ we have $\mathbb{E}[d(x^{j},x^{*})^{2}]\leq D^{2}$. Choosing $t_{k}=\frac{D}{L\sqrt{k+1}}$, $j=\lceil n/2\rceil$, and appealing to [7, Lemma 8.27(b)] gives
@@
The last statement of the theorem can be proven in exactly the same way as [7, Theorem 8.35(a)]. ∎
@@
The following lemma and its proof are straightforwardly adapted from [7, Lemma 8.39], which is itself an adaptation of the original [30, Lemma 2.1].
@@
With Lemma 6.3 in hand, the proof below is a rewrite of [7, Theorem 8.40] with the metric $d$ in place of the Euclidean distance.
@@
Applying [7, Lemma 8.27(b)] gives the bound
@@
and the final estimate follows. The last statement of the theorem can be proven in exactly the same way as [7, Theorem 8.40(a)]. ∎
@@ 7 Subgradient methods for the median problem
Lastly, as we have already noted, Busemann subgradient algorithms handle a nontrivial feasible region $C$ routinely, assuming that the projection $P_{C}$ is accessible. Solving an optimization problem of the form $\inf_{C}\sum_{i=1}^{m}f_{i}$ using the proximal methods of [5], on the other hand, forces a choice of approaches, all with disadvantages. One might model the hard constraint $x\in C$ using the indicator function $\delta_{C}$ as another summand, but the developments of [5, 31] ostensibly concern only Lipschitz components. An exact penalty function approach [14, Proposition 6.3] lets us replace $\delta_{C}$ by a multiple of the distance function $\mbox{dist}_{C}$, which is Lipschitz. However, even if all the components $f_{i}$ are $L$-Lipschitz, the required multiple must exceed $mL$, worsening the complexity estimates in comparison with the subgradient approach. Finally, we might replace proximal steps for each $f_{i}$ with those for $f_{i}+\delta_{C}$, but unlike the unconstrained case, these steps may be hard to compute even when each $f_{i}$ is a distance function.
@@ 8 Computational experiments
We consider the BHV tree space $\mathcal{T}_{n}$ of binary trees on $n$ labelled leaves [10]. There are (2n3)!!(2n-3)!! such binary trees, each with n2n-2 internal edges. The space $\mathcal{T}_{n}$ models all such binary trees by ascribing an (n2)(n-2)-dimensional orthant $0,\infty)^{n-2}$ to each tree so that a point in each orthant describes a particular binary tree topology with a prescribed choice of nonnegative internal edge lengths. Tree space has proved an interesting model for comparison and averaging of phylogenetic trees. The space $\mathcal{T}_{n}$ was shown in [10] to be CAT(0), is complete, and has the geodesic extension property, by [13, Proposition II.5.10] and [10, p. 743].
@@
Now we consider three trees in $\mathcal{T}_{4}$ whose median lies on the common boundary ray of three quadrants, with each quadrant containing one of the trees. We refer to this common boundary ray as the spine. A figure illustrating the local geometry of this setup can be found in [10, Figure 10]. The three neighboring quadrants arise by permuting the labels on the leaves of the left subtree of the leftmost tree in Figure 8.3. The spine corresponds to the second tree in this same figure, obtained by contracting the lower internal edge so that the left subtree leaves become siblings. The label $L$ on the separated leaf remains fixed throughout.
@@
The inequality in [13, Lemma II.8.21(2)] asserts
@@
By [13, Proposition II.8.22], we know that $r(\delta)$ is the unique minimizer of $b_{\xi}$ on the sphere of radius $\delta$ around $\bar{x}$, denoted by $S_{\delta}(\bar{x})$. Thus, $\min_{z\in S_{\delta}(\bar{x})}b_{\xi}(z)=b_{\xi}(r(\delta))=-\delta$. Likewise, it is easy to see from (27) that $\lim_{n\to\infty}b_{\xi}(r_{n}(\delta))=-\delta$. As the limit $\bar{z}=\lim_{n\to\infty}r_{n}(\delta)$ is in $S_{\delta}(\bar{x})$, we conclude that
@@
Conversely, let us suppose that $r_{n}(\delta)\to r(\delta)$ for all $\delta>0$ and prove $b_{\xi_{n}}\to b_{\xi}$ uniformly on bounded subsets. It suffices to prove $b_{\xi_{n}}\to b_{\xi}$ uniformly on each ball $B_{\delta}(\bar{x})$ for all $\delta>0$. Fix $\delta>0$ and let $\varepsilon>0$ be arbitrary. The inequality in [13, Lemma II.8.21(1)] asserts the existence of $R>0$ such that
@@
To prove the implication (ii) $\Rightarrow$ (i), we suppose that $f$ has a horospherically convex epigraph. Let $x\in X$ and consider the point $(x,f(x))$ in the boundary of $\operatorname{epi}f$. Since $\operatorname{epi}f$ is horospherically convex, there exists a ray $r\colon\mathbb{R}_{+}\to X\times\mathbb{R}$ issuing from $(x,f(x))$ such that $\operatorname{epi}f\subseteq b_{r}^{-1}((-\infty,0])$. The point at infinity $r(\infty)$ lies in $(X\times\mathbb{R})^{\infty}$, which is homeomorphic to the spherical join of $X^{\infty}$ and $\mathbb{R}^{\infty}$ (see [13, Example II.8.11(6)]), so we may assume $b_{r}(z,t)=\alpha b_{r_{1}}(z)+\beta b_{r_{2}}(t)$ for rays $r_{1}\colon\mathbb{R}_{+}\to X,r_{2}\colon\mathbb{R}_{+}\to\mathbb{R}$ and scalars $\alpha,\beta\geq 0$ satisfying $\alpha^{2}+\beta^{2}=1$. Realizing that $r_{2}$ must be of the form $r_{2}(t)=f(x)\pm t$, we derive the following inequality for all $(y,s)\in\operatorname{epi}f$:
@@
Thus $f$ is a supremum of nonnegatively scaled Busemann functions, whose epigraphs are easily seen to be horoballs in $X\times\mathbb{R}$. The epigraph of $f$ is therefore an intersection of horoballs, and such intersections are horospherically convex when $X$ is proper: see [16, Theorem 10] for a proof in the manifold case which generalizes, nearly verbatim, to proper Hadamard spaces. This proves (ii).
@@
Despite the initial requirement that this inequality holds for all $y\in X$, it is shown in [28, Remark 3.3(iv)] that the subdifferential of a geodesically convex function depends only on the function locally. Thus it suffices to verify (35) on a small ball around $\bar{x}$ contained in the quadrant $Q=\mathbb{R}_{+}\times\mathbb{R}_{+}\times\left\{0\right\}$, say $B_{1/5}(\bar{x})$. For $y=(y_{1},y_{2},0)\in Q$ we have

### v4

@@ 6 Subgradient-based minimization of sums
We denote the projection onto a nonempty closed geodesically convex set $C$ by $P_{C}$ (well-defined by [14, Proposition II.2.4]). Define an oracle Busemann that accepts a Busemann subdifferentiable function $g\colon C\to\mathbb{R}$ and a point $x\in C$, and returns a Busemann subgradient for $g$ at $x$: $\xi,s]\leftarrow\texttt{Busemann}(g,x)$. If $\xi,s]$ is a Busemann subgradient for $g$ at $x$, we will use the notation $r_{x,\xi}$ to denote the ray issuing from $x$ with direction $\xi$.
@@
    The function $f$ decomposes as $f=\sum_{i=1}^{m}f_{i}$, where each $f_{i},i=1,\dots,m$ is Busemann subdifferentiable on an open set containing $C$.
@@
$d(x^{+},y)^{2}$ $=d\left(P_{C}(r\hskip-1.42271pt\left(st\right)),P_{C}(y)\right)^{2}$
$\leq d\left(r\hskip-1.42271pt\left(st\right),y\right)^{2}$
@@
The second line uses nonexpansivity of the projection $P_{C}$, see [7, Theorem 2.1.12]. Letting $\delta\to\infty$ and using once more (5) we deduce
@@
The analysis of the stochastic subgradient method borrows from [6, Lemma 3.6] and [8, Theorem 8.35]. Define the notation $f^{k}_{\text{best}}:=\min_{i=1,\dots,k}f(x^{i})-f_{\text{opt}}$. For stochastic terminology, see [36].
@@
Finally, since the diameter of $C$ is at most $D$ we have $\mathbb{E}[d(x^{j},x^{*})^{2}]\leq D^{2}$. Choosing $t_{k}=\frac{D}{L\sqrt{k+1}}$, $j=\lceil n/2\rceil$, and appealing to [8, Lemma 8.27(b)] gives
@@
The last statement of the theorem can be proven in exactly the same way as [8, Theorem 8.35(a)]. ∎
@@
The following lemma and its proof are straightforwardly adapted from [8, Lemma 8.39], which is itself an adaptation of the original [32, Lemma 2.1].
@@
With Lemma 6.3 in hand, the proof below is a rewrite of [8, Theorem 8.40] with the metric $d$ in place of the Euclidean distance.
@@
Applying [8, Lemma 8.27(b)] gives the bound
@@
and the final estimate follows. The last statement of the theorem can be proven in exactly the same way as [8, Theorem 8.40(a)]. ∎
@@ 7 Subgradient methods for the median problem
Lastly, as we have already noted, Busemann subgradient algorithms handle a nontrivial feasible region $C$ routinely, assuming that the projection $P_{C}$ is accessible. Solving an optimization problem of the form $\inf_{C}\sum_{i=1}^{m}f_{i}$ using the proximal methods of [6], on the other hand, forces a choice of approaches, all with disadvantages. One might model the hard constraint $x\in C$ using the indicator function $\delta_{C}$ as another summand, but the developments of [6, 33] ostensibly concern only Lipschitz components. An exact penalty function approach [15, Proposition 6.3] lets us replace $\delta_{C}$ by a multiple of the distance function $\mbox{dist}_{C}$, which is Lipschitz. However, even if all the components $f_{i}$ are $L$-Lipschitz, the required multiple must exceed $mL$, worsening the complexity estimates in comparison with the subgradient approach. Finally, we might replace proximal steps for each $f_{i}$ with those for $f_{i}+\delta_{C}$, but unlike the unconstrained case, these steps may be hard to compute even when each $f_{i}$ is a distance function.
@@ 8 Computational experiments
We consider the BHV tree space $\mathcal{T}_{n}$ of binary trees on $n$ labelled leaves [11]. There are (2n3)!!(2n-3)!! such binary trees, each with n2n-2 internal edges. The space $\mathcal{T}_{n}$ models all such binary trees by ascribing an (n2)(n-2)-dimensional orthant $0,\infty)^{n-2}$ to each tree so that a point in each orthant describes a particular binary tree topology with a prescribed choice of nonnegative internal edge lengths. Tree space has proved an interesting model for comparison and averaging of phylogenetic trees. The space $\mathcal{T}_{n}$ was shown in [11] to be CAT(0), is complete, and has the geodesic extension property, by [14, Proposition II.5.10] and [11, p. 743].
@@
Now we consider three trees in $\mathcal{T}_{4}$ whose median lies on the common boundary ray of three quadrants, with each quadrant containing one of the trees. We refer to this common boundary ray as the spine. A figure illustrating the local geometry of this setup can be found in [11, Figure 10]. The three neighboring quadrants arise by permuting the labels on the leaves of the left subtree of the leftmost tree in Figure 8.3. The spine corresponds to the second tree in this same figure, obtained by contracting the lower internal edge so that the left subtree leaves become siblings. The label $L$ on the separated leaf remains fixed throughout.
@@
Acknowledgement: The authors thank two anonymous reviewers for many helpful and substantial suggestions that greatly improved this work.
@@
The inequality in [14, Lemma II.8.21(2)] asserts
@@
By [14, Proposition II.8.22], we know that $r(\delta)$ is the unique minimizer of $b_{\xi}$ on the sphere of radius $\delta$ around $\bar{x}$, denoted by $S_{\delta}(\bar{x})$. Thus, $\min_{z\in S_{\delta}(\bar{x})}b_{\xi}(z)=b_{\xi}(r(\delta))=-\delta$. Likewise, it is easy to see from (27) that $\lim_{n\to\infty}b_{\xi}(r_{n}(\delta))=-\delta$. As the limit $\bar{z}=\lim_{n\to\infty}r_{n}(\delta)$ is in $S_{\delta}(\bar{x})$, we conclude that
@@
Conversely, let us suppose that $r_{n}(\delta)\to r(\delta)$ for all $\delta>0$ and prove $b_{\xi_{n}}\to b_{\xi}$ uniformly on bounded subsets. It suffices to prove $b_{\xi_{n}}\to b_{\xi}$ uniformly on each ball $B_{\delta}(\bar{x})$ for all $\delta>0$. Fix $\delta>0$ and let $\varepsilon>0$ be arbitrary. The inequality in [14, Lemma II.8.21(1)] asserts the existence of $R>0$ such that
@@
To prove the implication (ii) $\Rightarrow$ (i), we suppose that $f$ has a horospherically convex epigraph. Let $x\in X$ and consider the point $(x,f(x))$ in the boundary of $\operatorname{epi}f$. Since $\operatorname{epi}f$ is horospherically convex, there exists a ray $r\colon\mathbb{R}_{+}\to X\times\mathbb{R}$ issuing from $(x,f(x))$ such that $\operatorname{epi}f\subseteq b_{r}^{-1}((-\infty,0])$. The point at infinity $r(\infty)$ lies in $(X\times\mathbb{R})^{\infty}$, which is homeomorphic to the spherical join of $X^{\infty}$ and $\mathbb{R}^{\infty}$ (see [14, Example II.8.11(6)]), so we may assume $b_{r}(z,t)=\alpha b_{r_{1}}(z)+\beta b_{r_{2}}(t)$ for rays $r_{1}\colon\mathbb{R}_{+}\to X,r_{2}\colon\mathbb{R}_{+}\to\mathbb{R}$ and scalars $\alpha,\beta\geq 0$ satisfying $\alpha^{2}+\beta^{2}=1$. Realizing that $r_{2}$ must be of the form $r_{2}(t)=f(x)\pm t$, we derive the following inequality for all $(y,s)\in\operatorname{epi}f$:
@@
Thus $f$ is a supremum of nonnegatively scaled Busemann functions, whose epigraphs are easily seen to be horoballs in $X\times\mathbb{R}$. The epigraph of $f$ is therefore an intersection of horoballs, and such intersections are horospherically convex when $X$ is proper: see [17, Theorem 10] for a proof in the manifold case which generalizes, nearly verbatim, to proper Hadamard spaces. This proves (ii).
@@
Despite the initial requirement that this inequality holds for all $y\in X$, it is shown in [30, Remark 3.3(iv)] that the subdifferential of a geodesically convex function depends only on the function locally. Thus it suffices to verify (35) on a small ball around $\bar{x}$ contained in the quadrant $Q=\mathbb{R}_{+}\times\mathbb{R}_{+}\times\left\{0\right\}$, say $B_{1/5}(\bar{x})$. For $y=(y_{1},y_{2},0)\in Q$ we have